# A Progressive Image Restoration Network for High-order Degradation Imaging in Remote Sensing

### v1

@@ IV-B5 DDLB architecture
TABLE VIII: Comparison of different SRx4 models on various scene classes of WHU-RS19 (PSNR ($dB$)/SSIM). The best and second-best results are highlighted in Bold font and underlined ones, respectively.
WHU-RS19 Airport Beach Bridge Commercial Desert Farmland FootballField Forest Industrial Meadow
Methods PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM
CTNet [48] 24.88/0.655 35.59/0.938 30.01/0.835 22.09/0.552 35.71/0.892 32.09/0.790 25.39/0.701 25.56/0.470 23.62/0.605 33.58/0.797
HSENet [49] 25.64/0.695 37.12/0.947 30.86/0.851 22.67/0.598 36.25/0.897 32.75/0.804 26.28/0.738 26.02/0.511 24.48/0.657 33.92/0.800
D-DBPN [45] 25.59/0.694 36.30/0.942 30.92/0.851 22.72/0.599 36.06/0.895 32.62/0.802 26.24/0.737 25.97/0.506 24.46/0.655 33.86/0.800
RDN [17] 25.52/0.693 36.85/0.945 30.85/0.852 22.62/0.595 36.21/0.897 32.72/0.803 26.15/0.738 25.93/0.509 24.36/0.652 33.89/0.800
TransENet [28] 25.69/0.696 37.40/0.948 30.88/0.854 22.73/0.600 36.03/0.897 32.74/0.804 26.39/0.745 26.08/0.517 24.54/0.659 33.92/0.801
RCAN [10] 25.68/0.702 37.07/0.946 30.93/0.855 22.76/0.606 36.20/0.897 32.86/0.806 26.35/0.746 26.06/0.516 24.61/0.665 33.90/0.800
SAN [20] 25.53/0.692 36.65/0.944 30.81/0.850 22.61/0.593 36.19/0.896 32.69/0.802 26.11/0.733 25.94/0.506 24.40/0.651 33.90/0.799
MHAN [24] 25.64/0.696 37.20/0.947 30.69/0.851 22.70/0.598 36.24/0.897 32.75/0.803 26.27/0.740 25.98/0.512 24.49/0.656 33.93/0.800
SwinIR [12] 25.79/0.704 37.49/0.949 30.99/0.855 22.81/0.609 36.31/0.898 32.90/0.807 26.47/0.750 26.12/0.521 24.66/0.668 33.96/0.801
HAUNet [27] 25.67/0.695 37.16/0.947 30.87/0.852 22.71/0.598 36.17/0.897 32.89/0.806 26.42/0.742 26.11/0.518 24.55/0.658 33.95/0.800
TTST [26] 25.76/0.703 37.21/0.948 30.85/0.854 22.80/0.607 36.23/0.898 32.89/0.806 26.46/0.748 26.09/0.519 24.68/0.667 33.93/0.801
CAT [46] 25.84/0.706 37.15/0.947 31.01/0.855 22.85/0.611 36.33/0.898 33.00/0.809 26.62/0.755 26.11/0.521 24.78/0.673 34.04/0.802
RGT [47] 25.93/0.710 37.60/0.949 31.25/0.858 22.91/0.613 36.37/0.898 33.05/0.810 26.72/0.758 26.15/0.523 24.85/0.675 34.06/0.802
DAT [22] 25.92/0.710 37.54/0.949 31.28/0.858 22.92/0.614 36.41/0.898 33.05/0.810 26.71/0.758 26.16/0.522 24.87/0.676 34.06/0.802
HDI-PRNet 26.10/0.712 36.88/0.948 31.61/0.862 23.05/0.615 36.40/0.898 33.09/0.811 26.87/0.757 26.19/0.526 24.94/0.674 34.07/0.803
WHU-RS19 Mountain Park Parking Pond Port RailwayStation Residential River Viaduct Average
Methods PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM
CTNet [48] 23.01/0.433 25.38/0.585 23.10/0.676 30.22/0.838 24.27/0.736 23.05/0.501 21.79/0.586 25.98/0.606 23.08/0.562 26.74/0.671
HSENet [49] 23.43/0.482 25.87/0.612 24.30/0.738 30.81/0.850 25.04/0.771 23.96/0.570 22.62/0.643 26.54/0.641 24.13/0.631 27.49/0.707
D-DBPN [45] 23.37/0.474 25.81/0.612 24.29/0.738 30.72/0.848 25.14/0.771 23.92/0.569 22.64/0.640 26.48/0.637 24.20/0.635 27.42/0.705
RDN [17] 23.39/0.479 25.83/0.612 24.16/0.733 30.74/0.848 25.01/0.769 23.96/0.572 22.52/0.637 26.47/0.638 24.16/0.635 27.42/0.705
TransENet [28] 23.48/0.488 25.92/0.615 24.52/0.752 30.83/0.851 25.20/0.775 24.05/0.572 22.77/0.651 26.57/0.644 24.23/0.638 27.56/0.710
RCAN [10] 23.46/0.489 25.94/0.617 24.39/0.752 30.82/0.852 25.14/0.777 24.10/0.583 22.68/0.652 26.55/0.644 24.37/0.647 27.55/0.713
SAN [20] 23.37/0.474 25.81/0.611 24.11/0.731 30.72/0.848 24.96/0.768 23.87/0.564 22.53/0.637 26.42/0.636 24.03/0.623 27.38/0.702
MHAN [24] 23.42/0.481 25.89/0.614 24.19/0.740 30.78/0.850 25.08/0.772 23.96/0.571 22.59/0.644 26.50/0.640 24.18/0.634 27.48/0.707
SwinIR [12] 23.50/0.492 25.95/0.619 24.75/0.762 30.87/0.853 25.29/0.780 24.15/0.583 22.75/0.657 26.62/0.648 24.36/0.648 27.65/0.715
HAUNet [27] 23.46/0.482 25.87/0.612 24.71/0.760 30.80/0.851 25.17/0.774 24.04/0.569 22.74/0.652 26.59/0.645 24.22/0.633 27.56/0.709
TTST [26] 23.50/0.491 25.94/0.617 24.67/0.756 30.90/0.853 25.28/0.779 24.09/0.580 22.73/0.654 26.61/0.647 24.34/0.646 27.61/0.714
CAT [46] 23.50/0.489 25.96/0.620 24.91/0.770 30.91/0.854 25.38/0.782 24.27/0.589 22.86/0.662 26.64/0.649 24.48/0.656 27.70/0.718
RGT [47] 23.54/0.496 26.06/0.622 25.01/0.772 30.99/0.854 25.45/0.785 24.33/0.594 22.94/0.664 26.70/0.651 24.56/0.658 27.79/0.720
DAT [22] 23.55/0.495 26.05/0.622 25.06/0.772 30.97/0.855 25.43/0.784 24.34/0.594 22.92/0.664 26.70/0.653 24.54/0.659 27.79/0.720
HDI-PRNet 23.62/0.498 26.21/0.623 25.29/0.778 31.09/0.854 25.58/0.783 24.52/0.601 23.21/0.668 26.74/0.652 24.60/0.657 27.87/0.721
@@
Figure 4: PSNR violin-plot of different x2 SR methods on WHU-RS19 dataset.
IV-C Comparisons with other methods on WHU-RS19 dataset
In this section, we assess the effectiveness of our proposed method by comparing it with popular image restoration methods using the WHU-RS19 dataset. Tables VI, VII, and VIII present the quantitative analysis of various evaluation metrics on different super-resolution scales. Our proposed HDI-PRNet achieves the best reconstruction performance in terms of average reconstruction metrics, showing average improvements of 0.35dB, 0.24dB, and 0.08dB. Furthermore, it also produces optimal or suboptimal reconstruction results in most land-use categories.
Fig. 3 shows a visualization comparison of the restoration results obtained by different methods in the WHU-RS19 dataset. Our method significantly alleviates the degradation caused by noise, blur, and downsampling, resulting in sharper images that are visually close to the ground truth. Although other methods have attained some success, they cannot fully restore the details of texture and structure. Specifically, the restoration results of CTNet and HSENet have many artifacts. Approaches such as DAT and HAUNet have complex designed architectures that can extract richer features, but they all produce over-smoothing results. In contrast, our method not only removes noise but also effectively restores image details and edges, obtaining the clearest and most accurate remote sensing images.
@@ IV-C Comparisons with other methods on WHU-RS19 dataset
Fig. 4 displays a violin-plot that serves as a tool for assessing the stability in performance of the high-order restoration framework. The plot includes three primary lines: the first quartile (Q1), the median, and the third quartile (Q3), which are useful for assessing the result’s dispersion. The Q1, median and Q3 values of the proposed method are the highest, indicating better overall performance compared to other approaches. The curve in the violin-plot represents a kernel density map, reflecting the data distribution shape. As depicted in Fig. 4, our method exhibits a more concentrated data distribution with less variance, contributing to greater model stability than alternative methods.
@@
TABLE IX: Comparison of different restoration models on real satellite images.
@@ IV-D Comparisons on real remote sensing images
Since real remote sensing images lack HR ground truth, we employ BRISQUE to evaluate the predicted results. As shown in Table IX, HDI-PRNet achieves the best or second-best results on real images from five different satellites. Although MHAN records the highest score on Jilin-1, its predictions are visually over-smoothed, which is not preferable. Overall, these significant improvements in both visual quality and reconstruction performance metrics strongly show the effectiveness and practical utility of HDI-PRNet to restore real remote sensing images, further highlighting the excellence of our proposed HDI-PRNet.
@@ V Conclusion
In this paper, we proposed a novel remote sensing high-order restoration framework. Our framework is based on the inverse design of the high-order degradation model to gradually restore the image, which provides a theoretical basis for the design of the DL framework. In addition, we introduced a denoising module that simulates the proximal mapping operator to learn the denoising prior representation of the current image. To solve the deblurring problem, we proposed a deblurring module that ensures the accurate representation of the degradation operator based on the expansion of the Neumann series and dual-domain degradation learning. We also introduced an intermediate loss in the middle of the network so that each module has supervision and the overall network is more interpretable. We performed several high-order degradation reconstruction experiments on WHU-RS19 and real remote sensing images. Experimental results show that our HDI-PRNet achieves competitive results on both synthetic and real images.

### v2

@@ IV-B5 DDLB architecture
TABLE IX: Comparison of different SR $\times$4 models on various scene classes of UCMerced (PSNR ($dB$)). The best and second-best results are highlighted in Bold font and underlined ones, respectively.
Class CTNet HSENet D-DBPN RDN TransE- RCAN SAN MHAN SwinIR HAUNet TTST CAT RGT DAT HDI-
[58] [59] [55] [22] Net[33] [14] [25] [29] [17] [32] [31] [56] [57] [27] PRNet
Agricultural 25.07 25.33 25.28 25.36 25.55 25.52 25.35 25.41 25.60 25.63 25.59 25.67 25.73 25.70 25.72
Airplane 24.91 25.76 25.74 25.63 25.76 25.86 25.64 25.75 25.91 25.94 25.86 25.96 26.03 26.02 26.33
Baseball Diamond 29.15 29.71 29.72 29.74 29.71 29.74 29.65 29.73 29.87 29.93 29.86 29.87 29.88 29.93 30.13
Beach 31.22 31.66 31.50 31.51 31.69 31.59 31.57 31.57 31.70 31.77 31.71 31.74 31.79 31.82 31.80
Buildings 22.79 23.68 23.75 23.57 23.74 23.80 23.54 23.60 23.82 23.88 23.83 23.97 23.99 24.08 24.24
Chaparral 22.93 23.73 23.65 23.72 23.94 23.76 23.61 23.71 23.93 23.97 23.95 23.91 23.79 23.89 23.62
Dense Residential 22.97 23.67 23.70 23.57 23.77 23.71 23.55 23.61 23.80 23.91 23.77 23.90 23.98 23.97 24.30
Forest 25.40 25.73 25.73 25.72 25.79 25.78 25.68 25.74 25.82 25.84 25.80 25.83 25.87 25.86 25.88
Freeway 25.23 26.43 26.46 26.44 26.47 26.63 26.30 26.48 26.52 26.67 26.52 26.69 26.79 26.82 27.01
Golfcourse 29.33 29.81 29.76 29.81 29.92 29.83 29.78 29.83 29.90 29.93 29.83 29.93 29.99 30.00 30.12
Harbor 20.49 21.26 21.32 21.14 21.39 21.25 21.12 21.18 21.44 21.65 21.34 21.59 21.57 21.67 22.05
Intersection 24.04 24.72 24.78 24.66 24.77 24.84 24.67 24.73 24.84 24.91 24.86 24.96 25.01 25.03 25.27
Medium Residential 23.31 23.88 23.90 23.84 23.96 23.93 23.81 23.87 24.02 24.07 23.98 24.08 24.13 24.19 24.29
Mobile Home Park 21.50 22.17 22.21 22.03 22.14 22.13 22.09 22.06 22.28 22.35 22.19 22.39 22.45 22.43 22.71
Overpass 23.88 24.82 24.90 24.86 24.89 25.02 24.69 24.89 25.02 25.04 24.95 25.11 25.25 25.27 25.46
Parking Lot 20.02 20.71 20.76 20.57 20.94 20.78 20.57 20.65 20.98 21.11 20.85 21.13 21.16 21.23 21.49
River 25.83 26.16 26.18 26.16 26.24 26.18 26.13 26.20 26.24 26.30 26.22 26.24 26.26 26.28 26.32
Runway 26.72 27.87 27.70 27.63 27.86 27.81 27.65 27.77 27.87 28.09 28.02 27.96 28.14 28.13 28.59
Sparse Residential 25.27 25.81 25.82 25.81 25.87 25.87 25.71 25.82 25.97 25.99 25.91 25.99 26.01 26.04 26.13
Storage Tanks 24.91 25.63 25.60 25.63 25.64 25.59 25.51 25.60 25.73 25.82 25.68 25.80 25.93 25.90 26.20
Tennis Court 25.52 26.15 26.16 26.11 26.25 26.21 26.03 26.14 26.32 26.38 26.24 26.34 26.44 26.48 26.71
TABLE X: Comparison of different SR $\times$3 models on various scene classes of NWPU-RESISC45 (PSNR ($dB$)). The best and second-best results are highlighted in Bold font and underlined ones, respectively.
Class CTNet HSENet D-DBPN RDN TransE- RCAN SAN MHAN SwinIR HAUNet TTST CAT RGT DAT HDI-
[58] [59] [55] [22] Net[33] [14] [25] [29] [17] [32] [31] [56] [57] [27] PRNet
Airplane 26.46 26.66 26.98 26.89 26.95 26.94 26.92 27.00 27.06 27.08 27.08 27.14 27.12 27.16 27.35
Airport 26.54 26.74 27.08 27.11 27.03 27.16 27.04 27.16 27.17 27.18 27.21 27.27 27.31 27.34 27.49
Baseball Diamond 27.67 27.89 28.23 28.22 28.20 28.29 28.18 28.28 28.32 28.33 28.29 28.38 28.44 28.45 28.63
Basketball Court 26.09 26.26 26.57 26.54 26.50 26.59 26.50 26.61 26.57 26.67 26.60 26.77 26.72 26.69 26.96
Beach 27.92 28.01 28.17 28.25 28.26 28.33 28.24 28.31 28.35 28.45 28.35 28.36 28.42 28.44 28.57
Bridge 29.13 29.46 29.81 29.80 29.77 29.81 29.78 29.83 29.87 29.99 29.93 29.98 30.06 30.03 30.31
Chaparral 23.80 23.88 24.18 24.06 24.15 24.21 23.98 24.16 24.16 24.27 24.15 24.20 24.27 24.25 24.45
Church 23.78 23.95 24.25 24.18 24.22 24.25 24.17 24.28 24.29 24.35 24.25 24.36 24.41 24.42 24.69
Circular Farmland 29.17 29.36 29.67 29.68 29.65 29.76 29.64 29.77 29.75 29.83 29.74 29.83 29.88 29.89 30.11
Cloud 29.27 29.30 29.60 30.37 30.36 30.41 30.10 30.20 30.56 30.70 30.49 30.44 30.62 30.57 30.83
Commercial Area 24.17 24.37 24.62 24.59 24.66 24.66 24.60 24.70 24.72 24.77 24.73 24.84 24.81 24.86 25.11
Dense Residential 22.94 23.09 23.35 23.29 23.39 23.33 23.31 23.42 23.41 23.50 23.41 23.50 23.47 23.56 23.78
Desert 30.13 30.21 30.36 30.43 30.39 30.45 30.35 30.45 30.54 30.60 30.47 30.58 30.56 30.59 30.64
Forest 26.49 26.56 26.69 26.69 26.72 26.72 26.67 26.75 26.76 26.80 26.74 26.78 26.79 26.80 26.87
Freeway 26.66 26.87 27.20 27.20 27.21 27.31 27.16 27.29 27.30 27.40 27.27 27.39 27.46 27.41 27.60
Golf Course 29.13 29.35 29.60 29.68 29.71 29.74 29.60 29.77 29.80 29.87 29.78 29.84 29.87 29.91 30.06
Ground Track Field 26.23 26.46 26.78 26.79 26.78 26.88 26.77 26.89 26.87 26.98 26.95 26.98 27.04 27.05 27.22
Harbor 21.73 21.91 22.35 22.21 22.34 22.20 22.18 22.31 22.36 22.42 22.39 22.46 22.48 22.50 22.89
Industrial Area 25.41 25.65 26.02 25.92 26.03 25.99 25.97 26.06 26.01 26.15 26.08 26.13 26.20 26.19 26.53
Intersection 23.30 23.48 23.82 23.77 23.79 23.85 23.75 23.89 23.86 23.90 23.92 23.94 23.99 24.00 24.27
Island 31.12 31.22 31.44 31.48 31.43 31.56 31.53 31.64 31.66 31.69 31.55 31.69 31.75 31.74 31.85
Lake 29.21 29.29 29.55 29.57 29.55 29.61 29.54 29.60 29.64 29.64 29.61 29.68 29.68 29.68 29.78
Meadow 31.08 31.09 31.25 31.28 31.35 31.32 31.25 31.30 31.32 31.39 31.31 31.41 31.42 31.41 31.51
Medium Residential 24.48 24.66 24.91 24.85 24.89 24.91 24.85 24.96 24.96 25.00 24.97 25.02 25.03 25.07 25.25
Mobile Home Park 23.10 23.29 23.64 23.55 23.63 23.56 23.58 23.66 23.64 23.68 23.72 23.82 23.73 23.79 24.12
Mountain 26.74 26.83 26.92 26.97 26.97 27.01 26.93 27.01 27.02 27.06 27.03 27.06 27.07 27.10 27.15
Overpass 25.16 25.42 25.87 25.85 25.84 25.93 25.79 25.93 25.91 26.01 25.98 26.01 26.09 26.09 26.36
Palace 23.72 23.91 24.21 24.15 24.19 24.26 24.16 24.29 24.28 24.32 24.27 24.35 24.38 24.40 24.58
Parking Lot 21.62 21.75 22.07 21.97 21.99 21.98 21.96 22.03 22.05 22.16 22.14 22.22 22.22 22.21 22.63
Railway 25.38 25.59 25.91 25.93 25.88 25.97 25.83 25.98 25.93 25.99 25.95 26.03 26.09 26.09 26.32
Railway Station 24.56 24.76 25.00 25.01 25.06 25.11 25.01 25.10 25.12 25.21 25.15 25.24 25.26 25.26 25.42
Rectangular Farmland 29.54 29.81 30.08 30.13 30.11 30.18 30.08 30.23 30.25 30.34 30.24 30.37 30.41 30.45 30.68
River 28.19 28.35 28.48 28.55 28.54 28.61 28.53 28.61 28.62 28.67 28.64 28.69 28.70 28.72 28.83
Roundabout 24.86 25.03 25.31 25.29 25.30 25.37 25.27 25.39 25.37 25.44 25.40 25.43 25.47 25.47 25.66
Runway 29.18 29.47 29.88 29.96 29.93 30.15 29.91 30.18 30.13 30.32 30.19 30.33 30.47 30.41 30.76
Sea Ice 26.61 27.01 27.39 27.25 27.39 27.29 27.06 27.27 27.25 27.55 27.41 27.50 27.48 27.61 28.00
Ship 26.59 26.77 27.18 27.10 27.12 27.16 27.11 27.19 27.19 27.21 27.23 27.27 27.31 27.30 27.46
Snowberg 22.71 22.78 23.14 23.08 23.12 23.14 23.04 23.18 23.18 23.22 23.22 23.24 23.24 23.29 23.51
Sparse Residential 26.41 26.56 26.85 26.82 26.83 26.86 26.80 26.89 26.89 26.90 26.91 26.93 26.97 26.96 27.11
Stadium 25.26 25.49 25.78 25.73 25.75 25.81 25.75 25.83 25.86 25.93 25.94 25.94 25.98 26.00 26.36
Storage Tank 24.19 24.36 24.82 24.72 24.77 24.78 24.71 24.83 24.77 24.93 24.85 24.94 24.98 25.00 25.33
Tennis Court 25.67 25.86 26.11 26.10 26.10 26.15 26.06 26.19 26.17 26.24 26.20 26.28 26.27 26.28 26.53
Terrace 28.53 28.70 29.02 29.06 29.07 29.11 29.00 29.12 29.12 29.20 29.14 29.22 29.22 29.27 29.32
Thermal Power Station 26.21 26.39 26.74 26.75 26.82 26.86 26.74 26.86 26.85 26.91 26.89 26.93 27.00 26.99 27.23
Wetland 28.45 28.56 28.77 28.81 28.85 28.85 28.76 28.82 28.88 28.92 28.82 28.90 28.89 28.90 28.98
@@
Figure 4: PSNR violin-plot of different $\times$2 SR methods on WHU-RS19 dataset.
IV-C Comparisons with other methods on simulated datasets
In this section, we evaluate the superiority of our proposed method by comparing it with state-of-the-art image restoration methods. Table VI shows the quantitative analysis of the indicators of different algorithms in the ×2,×3,×4$\times$ 2$\times$ 3$\times$ 4× 2 , × 3 , × 4 image restoration task in the WHU-RS19, DOTA, UCMerced, RSSCN7 and NWPU-RESISC45 datasets. Table VI demonstrates that HDI-PRNet consistently achieves superior performance across all dataset tasks, achieving the highest average values of PSNR and SSIM. In particular, the proposed method shows significant improvements in reconstruction quality compared to state-of-the-art approaches, including DAT and TTST.
Tables VII, VIII, IX, and X present the PSNR results on the WHU-RS9, RSSCN7, UCMerced, and NWPU-RESISC45 datasets, demonstrating the comparative reconstruction performance in remote sensing image restoration tasks under varying sensors and land cover classes. The results reveal that HDI-PRNet achieves significantly superior reconstruction quality compared to TTST (a specialized method for remote sensing image reconstruction) and DAT (a general image super-resolution algorithm). In particular, HDI-PRNet exhibits consistent advantages across most land classes, further validating its geesubmit-neralizability and robustness in various remote sensing scenarios. These results highlight the method’s ability to deliver high-fidelity reconstructions while adapting to varying imaging conditions and scene complexities.
Fig. 3 presents a qualitative comparison of the restoration results in different methods. Our approach significantly mitigates image degradation caused by noise, blur, and downsampling, producing sharper and more visually realistic reconstructions. Although existing methods demonstrate certain improvements, they still struggle to fully recover fine textures and structural details. For instance, CTNet, HSENet, and TransENet exhibit noticeable artifacts and over-smoothing in their outputs. Although more sophisticated architectures such as DAT and HAUNet leverage richer feature representations, they fail to reconstruct edge details with sufficient clarity. In contrast, our method not only suppresses noise effectively, but also preserves and enhances critical image details and edges, yielding the clearest and most accurate remote sensing reconstructions among all compared approaches.
@@ IV-C Comparisons with other methods on simulated datasets
Fig. 4 shows a violin plot that serves as a tool to evaluate the performance stability of the HDI-PRNet restoration framework. The plot includes three primary lines: the first quartile (Q1), the median, and the third quartile (Q3), which are useful for assessing the result’s dispersion. The Q1, median and Q3 values of the proposed method are the highest, indicating better overall performance compared to other approaches. As depicted in Fig. 4, our method exhibits better reconstruction performance and can handle different data more stably.
@@
TABLE XI: Comparison of different restoration models on real satellite images.
@@ IV-D Comparisons on real remote sensing images
Since real remote sensing images lack HR ground truth, we employ BRISQUE to evaluate the predicted results. As shown in Table XI, HDI-PRNet achieves the best or second-best results on real images from five different satellites. Although MHAN records the highest score on Jilin-1, its predictions are visually over-smoothed, which is not preferable. Overall, these significant improvements in both visual quality and reconstruction performance metrics strongly show the effectiveness and practical utility of HDI-PRNet to restore real remote sensing images, further highlighting the excellence of our proposed HDI-PRNet.
@@ V Conclusion
In this paper, we proposed a novel remote sensing high-order restoration framework. Our framework is based on the inverse Markov process design of the high-order degradation model to gradually restore the image, which provides a theoretical basis for the design of the DL framework. In addition, we introduced a denoising module that simulates the proximal mapping operator to learn the denoising prior representation of the current image. To solve the deblurring problem, we proposed a deblurring module that ensures the accurate representation of the degradation operator based on the expansion of the Neumann series and dual-domain degradation learning. We also introduced an intermediate loss in the middle of the network so that each module has supervision and the overall network is more interpretable. We performed several high-order degradation reconstruction experiments on WHU-RS19 and real remote sensing images. Experimental results show that our HDI-PRNet achieves competitive results on both synthetic and real images.
Although HDI-PRNet has achieved excellent performance in the task of real remote sensing image restoration, there are still some limitations. To verify the effectiveness of the high-order model design, this paper only discussed the three classic image restoration tasks of denoising, deblurring, and super-resolution, but the real remote sensing imaging process also involves various other types of degradation, such as clouds and artifacts. The high-order degradation model composed of more types of degradation makes the restoration process of real images complex and challenging.